# Learning Diverse Robot Striking Motions with Diffusion Models and Kinematically Constrained Gradient Guidance

### v1

@@ V-B Sampling Procedure
We describe our guidance procedure in Algorithm 1. We predict a batch of trajectories $\hat{\tau}_{0}$ with our network $s_{\theta}(\tau_{i},i)$ (Lines 4 - 5). An intermediate sample $\tau_{i-1}^{\prime}$ is computed using the standard DDPM procedure (Lines 6 - 7). Then, the sample is updated by KCGG in Line 8. The function $F$ represents the forward kinematics chain for the robot arm. We define our constraint function as $c(\tau)=\min_{t\in[t_{s},t_{e}],q_{t}\in\tau}\|F(q_{t})-b_{t}\|_{2}^{2}$, which minimizes the difference between the robot end-effector and ball prediction location at certain points in time defined by the window, $[t_{s},t_{e}]$. Finally, we select a single best trajectory $\tau_{0}^{*}$ based on the computed norm $N$ from the final denoising step (Line 10). This allows us to utilize the computational efficiency of batched operations on the GPU and select the trajectory that best matches the given end-effector constraint. We refer to this step as “Batch Filter” and ablate this feature.
@@ VI-A Agile Robotic Domains
To evaluate our method, we perform experiments on two distinct domains: AirHockey and Table-Tennis.
@@ VI-A1 Simulated Domain (AirHockey)
We evaluate the air hockey simulation domain [6] on two tasks, Hit and Defend. In Hit, the objective is to score a goal with the desired shot type, Straight or Angled. If a goal is scored with the shot type behavior, it is considered a Success. In Defend, the goal is to block an incoming puck. Block is defined as when the robot’s end-effector successfully makes contact with the puck. Models are trained on 200 (100 per shot type) demonstrations in Hit and 100 demonstrations in Defend.
@@
Conditional Diffusion Model 98.3% 82.4% 90.4%
TABLE I: AirHockey Hit Results: Comparison of Success rates on Straight and Angled shot types.
VI-A2 Real-World Domain (Table Tennis)
A 7-DOF Barrett WAM system from [30] with perception improvements [50] is used to return table tennis serves from a ball launcher. We collect diverse training data through kinesthetic teaching and record four types of shots: Push, Lob, Slice, and Topspin. Balls are launched to two striking locations on the table: Center and Left. Models are trained on 120 demonstrations (15 per condition) and evaluated on three metrics: Success, Hit, and Correct Stroke rate. Success is defined as a ball return within the spin and height distribution of the demonstrated shot type’s data. A Hit is recorded when the ball contacts the end-effector, but fails the Success criteria. Finally, Correct Stroke is a sampled stroke that matches the desired shot type’s joint trajectory distribution and striking location.
VI-B Baselines
We benchmarked against the following LfD baselines: Behavioral cloning (BC) policies with Gaussian Mixture Models (GMM) [49], Probabilistic Movement Primitives (ProMP) [26], and Action Chunking with Transformers (ACT) [19]. We also compare against five diffusion baselines: Diffuser [23], Motion Planning Diffusion (MPD1) [24], Diffusion Policy [51], BEbhavior generation with ScOre-based Diffusion Policies (BESO) [41] and Movement Primitive Diffusion (MPD2) [42].
@@ VI-B Baselines
Unconstrained w/o Batch Filter 12.9% 0.159 20 7.95
Unconstrained 62.6% 0.161 20 8.05
@@
Unconstrained w/o Batch Filter 46.2% 0.162 20 8.10
Unconstrained 46.6% 0.163 20 8.15
@@
TABLE II: Air-Hockey Defend Results: We compare our constraint sampling method against imitation learning and diffusion baselines. “T” is the number of diffusion timesteps.
@@ VI-C Simulation Results: AirHockey
In this section, we discuss results on the simulated AirHockey Hit and Defend tasks. For each task, we test on 1000 initial start configurations. Our results for Hit (Table I) show that the diffusion model captures diverse motions best, achieving highest success rates for each shot type. Results for the Defend task (Table II), show KCGG enables unconditioned diffusion models to outperform all baselines, increasing hit rate by 18.9% over the best baseline. We use a fixed wall-clock time budget of 200 ms and tune sampling steps accordingly. Interestingly, KCGG does not improve puck-conditioned models. We hypothesize the conditioned base model overfits the score function to puck observations.
@@
Figure 3 shows a comparison of the joint angle trajectories generated by KCGG and the original constraint guidance, revealing KCGG more effectively achieves the desired end-effector trajectory while adhering to the data distribution. The trajectory produced by the original constraint guidance exhibits significant inconsistencies in the constrained region whereas KCGG samples a smooth trajectory.
VI-D Sampling Time Ablation
We perform a sensitivity analysis comparing against other diffusion baselines by varying the number of diffusion timesteps (Figure 4). Despite requiring more time per step, KCGG outperforms diffusion baselines across all sampling time budgets, except for MPD2 with sampling time budget less than $70$ ms. Wall-clock time is evaluated as sampling step time vary for each method. All experiments were run on a single Nvidia GTX 1080.
Figure 4: Sampling Speed vs Performance: We compare the Block Rate in Air-Hockey Defend across sampling times. KCGG scales to better performance with sampling time.
As the number of diffusion timesteps increases, the performance of the baseline methods plateau. In contrast, KCGG scales effectively, reaching a block rate nearing 85%, while a method such as MPD2 does not, despite performing better with few sampling steps. Additional denoising steps allows KCGG to better adhere the trajectory samples to both the training data distribution and task constraints.
@@ VI-D Sampling Time Ablation
TABLE III: Table Tennis Results: We compare the hit rates (H), success rates (S), and correct stroke rates (CS) for four different shot types across two different ball launcher configurations (left and center launches).
@@ VI-E Real-world Results: Table Tennis
We report results comparing diffusion models to imitation learning baselines in Table III with 50 samples per task. Expert demonstrations are used instead of traditional heuristics, which lack the flexibility to generate diverse table-tennis stroke profiles and are hard to design. Based on Air-Hockey Defend findings, we conditioned our models only on stroke type to avoid over-parameterizing the score function.
In terms of task performance, KCGG shows an average 17.3% improvement in success rate across all tasks, outperforming or matching the highest baseline success rates in six of eight tasks. Specifically, KCGG surpasses prior diffusion constraint guidance methods (MPD1) with a notable 124.5% improvement. We also note a modest average 4% improvement in hit rate over the baselines, attributed mainly to high hit rates in the push and lob strokes.
@@
We evaluated KCGG’s ability to express multimodal behaviors across four stroke types and two ball launch configurations. KCGG consistently demonstrated the desired behavior in each task, shown by high correct stroke rates. Conversely, ACT, despite conditioning on stroke style, struggled to generate correct left/center strokes from ball observations, yielding a lower correct stroke rate. For BESO and MPD2, we observed that conditioning on ball observation history caused overfitting and collapse into certain stroke modalities (e.g. BESO mostly samples Center Lob strokes), illustrating the importance of constraint-guidance. We created stronger variants of these baselines by replacing ball history conditioning with ”ball oracle” conditioning, a categorical label indicating Left or Center launch distribution from the demonstration. This provides direct information on the ball launch location, which significantly improved BESO and MPD2 performance, removing mode collapse for a stronger comparison. However, KCGG still outperformed these variants on average in hit and success rate. ProMP, which was not designed to capture multimodal behavior, was trained on data specific to each task, thus we exclude it from correct stroke rate evaluations. KCGG, however, effectively modeled trajectories across all conditions, accurately selecting the appropriate stroke based on the constraints.
@@ VIII Conclusion
Our work demonstrates KCGG’s effectiveness on small expert datasets in agile robotic tasks. Our algorithm is validated in both a simulated air hockey domain and on a real table tennis robot. Our real robot experiments show that we can learn diverse policies without access to a high-fidelity simulator. These contributions address critical issues such as sample inefficiency and performance in dynamic settings, promising broader applications in real-world robotics.

### v2

@@ V-B Sampling Procedure
We describe our guidance procedure in Algorithm 1. We predict a batch of trajectories, $\hat{\tau}_{0}$, with our network, $s_{\theta}(\tau_{i},i)$ (Lines 4 - 5). An intermediate sample, $\tau_{i-1}^{\prime}$, is computed using the standard DDPM procedure (Lines 6 - 7). Then, the sample is updated by KCGG in Line 8. The function $F$ represents the forward kinematics chain for the robot arm. We define our constraint function as $c(\tau)=\min_{t\in[t_{s},t_{e}],q_{t}\in\tau}\|F(q_{t})-b_{t}\|_{2}^{2}$, which minimizes the difference between the robot end-effector and ball prediction location at certain points in time defined by the window, $[t_{s},t_{e}]$. Finally, we select a single best trajectory, $\tau_{0}^{*}$, based on the computed norm, $N$, from the final denoising step (Line 10). This allows us to utilize the computational efficiency of batched operations on the GPU and select the trajectory that best matches the given end-effector constraint. We refer to this step as “Batch Filter” and ablate this feature.
Hyperparameter Value
Learning Rate 2e-4
Planning Horizon 100
Batch Size 32
Number of Training Diffusion Steps 100
Total Training Steps 10e4
Optimizer Adam
U-Net Convolutional Blocks (1, 4, 8)
TABLE I: Architecture Hyperparameters in both domains
@@ VI-A Agile Robotic Domains
We evaluate on two domains: AirHockey and Table Tennis. In each domain, the object pose is estimated by a Kalman Filter [49]. A predicted path of the object is obtained from a linear dynamics rollout in AirHockey and a ball dynamics model [50] in Table Tennis. We form the time window, $[t_{s},t_{e}]$, around a predetermined hit plane derived from the mean location of object hits in the collected demonstrations.
@@ VI-A1 Simulated Domain (AirHockey)
We evaluate the air hockey simulation domain [6] on two tasks, Hit and Defend. In Hit, the objective is to score a goal with the desired shot type, Straight or Angled. A goal scored with the desired shot type is considered a Success. In Defend, the goal is to block an incoming puck that is initialized with random position and velocity. A Block is defined as when the robot’s end-effector successfully makes contact with the puck. Models are trained on 200 (100 per shot type) demonstrations in Hit and 100 demonstrations in Defend. Demonstrations are collected from a heuristic motion planner.
VI-A2 Real-World Domain (Table Tennis)
A 7-DOF Barrett WAM system from [30] with perception improvements [51] is used to return serves from a ball launcher. Training data is collected through kinesthetic teaching, with four types of shots: Push, Lob, Slice, and Topspin. Balls are launched to two locations on the table: Center and Left, with natural ball variance from the launcher. Models are trained on 120 demonstrations (15 per condition) and evaluated on three metrics: Success, Hit, and Correct Stroke rate. Success is defined as a ball return within the spin and height distribution of the shot type’s demonstrations. A Hit is recorded when the ball contacts the end-effector, but fails the Success criteria. Finally, Correct Stroke is a sample that matches the desired shot type’s joint trajectory distribution and striking location.
@@ VI-A2 Real-World Domain (Table Tennis)
Diffusion (Base) 98.3% 82.4% 90.4%
TABLE II: AirHockey Hit Results: Comparison of Success rates on Straight and Angled shot types.
VI-B Experiment Details and Baselines
Due to the tight timing requirements of our domains, we use a mixture of both [23] and [53] as our baseline diffusion model, which we refer to as “Diffusion (Base)”. We utilize the 1D Convolutional Temporal U-Net from [23] due to its faster sampling speed over attention-based models. For conditioned models, we incorporate Feature-Wise Linear Modulation (FiLM) Layers [54] to modulate the convolutional features within the U-Net, as suggested by [53]. Conditioned models use state-based observations of joint positions, stroke type, and the last two puck or last ten ball observations in the AirHockey and Table Tennis domains, respectively. We follow the standard training procedure from DDPM [35] to train all diffusion baselines. Model hyperparameters used for all domains are shown in Table I.
We benchmark against the following LfD baselines: Behavioral cloning with Gaussian Mixture Models (BC-GMM) [52], Probabilistic Movement Primitives (ProMP) [26], and Action Chunking with Transformers (ACT) [19]. We also compare against four diffusion baselines: Diffusion (Base), Motion Planning Diffusion (MPD1) [24], BEbhavior generation with ScOre-based Diffusion Policies (BESO) [41] and Movement Primitive Diffusion (MPD2) [42].
@@ VI-B Experiment Details and Baselines
Diffusion (Base) 12.9% 0.159 20 7.95
Diffusion (Base) w/ Batch Filter 62.6% 0.161 20 8.05
@@
Diffusion (Base) 46.2% 0.162 20 8.10
Diffusion (Base) w/ Batch Filter 46.6% 0.163 20 8.15
@@
TABLE III: AirHockey Defend Results: We compare our constraint sampling method against imitation learning and diffusion baselines. “T” is the number of diffusion timesteps.
@@ VI-C Simulation Results: AirHockey
In this section, we discuss results on the simulated AirHockey Hit and Defend tasks. For each task, we test on 1000 initial start configurations. Our results for Hit (Table II) show that Diffusion (Base) captures diverse motions best, achieving highest success rates for each shot type. Results for the Defend task (Table III), show KCGG enables unconditioned diffusion models to outperform all baselines, increasing block rate by 18.9% over the best baseline. We use a fixed wall-clock time budget of 200 ms and tune sampling steps accordingly. Interestingly, KCGG does not improve puck-conditioned models. We hypothesize that the conditioned model overfits the score function to puck observations.
We conduct an experiment on Defend by using an obstacle avoidance constraint with KCGG instead, i.e. $c(\tau)=\min_{t\in[t_{s},t_{e}],q_{t}\in\tau}\frac{1}{\|F(q_{t})-b_{t}\|_{2}^{2}}$, and achieve a $2.9\%$ block rate, showing KCGG is compatible with other constraints.
@@
Constraint wrt. $\tau_{i-1}$ Constraint wrt. $\tau_{0}$
Gradient wrt. $\tau_{i-1}$ 63.5% (MPD1 [24]) 65.0%
Gradient wrt. $\tau_{i}$ 70.0% 85.2% (Ours)
TABLE IV: Success Rates reported ablating for design choices in KCGG in the AirHockey Defend domain
Figure 3 shows a comparison of a joint angle trajectory generated by KCGG and MPD1 [24], revealing that the trajectory produced by MPD1 [24] exhibits significant inconsistencies in the constrained region whereas KCGG samples a smooth trajectory from the data distribution.
Figure 4: Sampling Speed vs Performance: We compare the Block Rate in AirHockey Defend across sampling times. KCGG scales to better performance with sampling time.
@@
TABLE V: Table Tennis Results: We compare the hit rates (H), success rates (S), and correct stroke rates (CS) for four different shot types across two different ball launcher configurations (left and center launches) with 50 trials per condition.
VI-D Ablation Studies
Table IV reports ablation results for: (1) Calculating the gradient with respect to $\tau_{i-1}$ or $\tau_{0}$ and (2) Computing the constraint function with $\tau_{i-1}$ or $\tau_{0}$. We find that taking the gradient with respect to $\tau_{i}$ brings a small improvement of 6.5% over MPD1 and taking the constraint with respect to $\tau_{0}$ also provides a small 1.5% improvement. With both, our KCGG method achieves large gains over the baselines.
We perform a sensitivity analysis against other diffusion baselines by varying the number of diffusion timesteps (Figure 4). Wall-clock time is evaluated as sampling step time vary for each method. All experiments were run on a single Nvidia GTX 1080. Despite requiring more time per step, KCGG outperforms diffusion baselines across all sampling time budgets, except for MPD2 with sampling time budget less than $70$ ms. As the number of diffusion timesteps increases, the performance of the baseline methods plateau. In contrast, KCGG scales, reaching a block rate near 85%, while a method such as MPD2 does not, despite performing better with few sampling steps. Additional denoising steps allows KCGG to better adhere the trajectory samples to both the training data distribution and task constraints.
@@ VI-E Real-world Results: Table Tennis
Results comparing KCGG to all other baselines are reported in Table V (50 samples per task). Based on AirHockey findings, our models are conditioned only on the stroke type to avoid over-parameterizing the score function.
In terms of task performance, KCGG shows an average 17.3% improvement in success rate across all tasks, outperforming or matching the highest baseline success rates in six of eight tasks. Specifically, KCGG surpasses prior diffusion constraint guidance methods (MPD1) with a notable 124.5% improvement. We also note a modest average 4% improvement in hit rate over the baselines.
@@
We evaluated KCGG’s ability to express multimodal behaviors across four stroke types and two ball launch configurations. KCGG consistently demonstrated the desired behavior in each task, shown by high correct stroke rates. Conversely, ACT, despite conditioning on stroke style, struggled to generate the correct left/center strokes from ball observations, yielding a lower correct stroke rate. For BESO and MPD2, we observed that conditioning on ball observation history caused overfitting and collapse into certain stroke modalities (e.g. BESO mostly samples Center Lob strokes), illustrating the importance of constraint guidance. We created stronger variants of these baselines by replacing ball history conditioning with “ball oracle” conditioning, a categorical label indicating Left or Center launch distribution from the demonstration. This provides direct information on the ball launch location, which significantly improved BESO and MPD2 performance. However, KCGG still outperformed these variants on average in hit and success rate. ProMP, which was not designed to capture multimodal behavior, was trained on data specific to each task, thus we exclude it from correct stroke rate evaluations. KCGG, however, effectively modeled trajectories across all conditions, accurately selecting the appropriate stroke based on the constraints.
@@ VIII Conclusion
Our work demonstrates KCGG’s effectiveness on small expert datasets in agile robotic tasks, through evaluations in both a simulated air hockey domain and on a real table tennis robot. Our real robot experiments show that we can learn diverse policies without access to a high-fidelity simulator. These contributions address critical issues such as sample inefficiency and performance in dynamic settings, promising broader applications in real-world robotics.